# Listening to the Universe with Next Generation Ground-Based Gravitational-Wave Detectors

### v2

@@ Abstract
In this study, we use simple performance metrics to assess the science capabilities of future ground-based gravitational-wave detector networks—composed of A+ or Voyager upgrades to the LIGO, Virgo, and KAGRA observatories and proposed next generation observatories such as Cosmic Explorer and Einstein Telescope. These metrics refer to coalescences of binary neutron stars (BNSs) and binary black holes (BBHs) and include: (i) network detection efficiency and detection rate of cosmological sources as a function of redshift, (ii) signal-to-noise ratios and the accuracy with which intrinsic and extrinsic parameters would be measured, and (iii) enabling multimessenger astronomy with gravitational waves by accurate 3D localization and early warning alerts. We further discuss the science enabled by the small population of rare and extremely loud events. While imminent upgrades will provide impressive advances in all these metrics, next generation observatories will deliver an improvement of an order-of-magnitude or more in most metrics. In fact, a network containing two or three such facilities will detect half of all the BNS and BBH mergers up to a redshift of $z=1$ and $z=20$, respectively, give access to hundreds of BNSs and ten thousand BBHs with signal-to-noise ratios exceeding 100, readily localize hundreds to thousands of mergers to within $1\,{\rm deg^{2}}$ on the sky and better than 10% in luminosity distance, respectively, and consequently, enable mutlimessenger astronomy through follow-up surveys in the electromagnetic spectrum several times a week. Such networks will further shed light on potential cosmological merger populations and detect an abundance of high-fidelity BNS and BBH signals which will allow investigations of the high-density regime of matter at an unprecedented level and enable precision tests of general relativity in the strong-filed regime, respectively.
@@ I.3 Future of gravitational-wave astronomy
There have been extensive studies on the capability of next generation detectors and the scientific discoveries they enable Sathyaprakash:2009xt ; Punturo:2010zza ; Zhao:2010sz ; Sathyaprakash:2012jk ; Vitale:2016icu ; Vitale:2018nif ; Adhikari:2019zpy ; Hall:2019xmm ; Sathyaprakash:2019nnu ; Sathyaprakash:2019rom ; Sathyaprakash:2019yqt ; Kalogera:2019sui ; Maggiore:2019uih ; Evans:2021gyd ; Saleem:2021iwi ; Bailes:2021tot ; Kalogera:2021bya ; Yang:2021qge ; Jin:2022qnj ; Jin:2020hmc ; Jin:2021pcv ; Zhang:2019loq In this study we will focus in detail on the potential of the Voyager, ET, and CE detector generations using the Voy, ET, and CE noise curves shown in Fig. 1.
@@ I.4 Organization of the paper
In Sec. II we will introduce the detector networks we are considering in this study, the properties of the source populations we are examining, and the Fisher-matrix package that was developed to assess the measurement capabilities of detector networks. In the rest of the paper we will introduce and discuss the metrics used to assess the relative performance of different networks. In Sec. III we will discuss the efficiency of detector networks in detecting signals as a function of redshift (or luminosity distance). Detection efficiency will inform us of the completeness of the observed population with respect to the full underlying population of mergers. Sec. III will also discuss the cosmic merger rate of compact binary coalescences inferred using current LIGO-Virgo observations, a model for the redshift evolution of the star formation rate and metallicity, and the resulting detection rates of the networks as functions of the redshift. In Sec. IV we will present the evaluation of the chosen detector networks and a forecast of the science questions that could be addressed by different generations of networks. Enabling multimessenger astronomy will require accurate 3D sky localization of compact binary coalescences, which will be the focus of Sec. V. This Section will also include forecasts for early warning alerts that would be possible with future detectors, with alerts sent to astronomers 10 minutes, 5 minutes, and 2 minutes before the merger occurs and the corresponding 3D sky localization of the event. Sec. VI is devoted to discussing the science enabled by rare events with SNRs larger than 300 or even 1000, e.g. observing higher multipole modes, testing the black hole no-hair theorem to exquisite precession, etc. Sec. VII provides a brief summary, limitations of the study, and how they might be improved and other studies that would be necessary to firm up the science potential of future detectors.
@@ II.1 Detector networks
Although the choice of when a specific A+ configuration is included in a NG network seems arbitrary, this resulted from our judgement that when a NG detector comes online the corresponding region’s 2G+ detector(s) may no longer be in operation. For example, Virgo is not likely to be operational when ET is built, LIGO-Hanford and LIGO-Livingston may not be operating at the same time as CE. Should it prove necessary to assess the performance of a network comprised of a different combination of detectors, e.g. HLI+C, it would be straightforward to do so with gwbench.
@@ II.2 Source populations
Throughout this paper, we examine two source populations comprising either BNSs or BBHs. Each population was uniformly distributed in six redshift bins, $z\in[0.02,0.5],\,[0.5,1],\,[1,2],\,[2,4],\,[4,10],\,[10,50]$, with 250,000 injections per bin. The corresponding luminosity distances $D_{L}$ were calculated with the Planck18 cosmology of astropy. The injections were further uniformly sampled over sky positions ($\alpha$ and $\cos(\delta)$) and binary orientation angles ($\cos(\iota)$ and $\psi$), with right ascension $\alpha$, declination $\delta$, inclination $\iota$, and polarization angle $\psi$. The injected spins were chosen to be aligned with the orbital angular momentum ($\chi_{1x}=\chi_{1y}=\chi_{2x}=\chi_{2y}=0$) while the $z$-components were uniformly sampled, for BNS $\chi_{1z},\,\chi_{2z}\in[-0.05,0,05]$ and for BBH $\chi_{1z},\,\chi_{2z}\in[-0.75,0,75]$. The BNS masses were chosen to be normally distributed in $[1,2]\,M_{\odot}$ with mean $\mu=1.35\,M_{\odot}$ and standard deviation $\sigma=0.15\,M_{\odot}$. The BBH masses were chosen to follow the POWER+PEAK distribution described in the Second Gravitational-Wave Transient Catalog population paper Abbott:2020gyp , but with the secondary mass sampled uniformly in $[5\,M_{\odot},m_{1}]$. These parameters are summarized in Tab. 2. Note that differences in the simulated populations and network configurations require cautious comparison with other literature.
@@ II.3 Analysis methodology
We assess the performance of the chosen GW networks using a fast, Fisher-information Cutler:1994ys ; Poisson:1995ef ; Balasubramanian:1995bm ; Arun:2004hn based Python package, gwbench, that we developed and made publicly available Borhanian:2020ypi ; Appendix A reports the settings used in gwbench. The Fisher information formalism approximates the parameter posteriors to be Gaussian (assuming Gaussian noise), provides an analytic recipe to calculate the covariance matrix of the posteriors, and thus allows the computation of measurement error estimates for the gravitational waveform parameters. The implementation in gwbench incorporates the effect of Earth’s rotation, which is important for signals lasting for longer than about an hour, and numerical differentiation schemes, enabling access to a host of gravitational waveform models in LAL, the LSC Algorithm Library lalsuite —ultimately enabling the use of more sophisticated waveform models and, therefore, more accurate estimates. gwbench packs this in an easy-to-use fashion and further gives access to signal-to-noise ratio (SNR) calculations, various detector locations and noise curves, and basic sampling methods.
@@ III.2 Network efficiency
Figure 2: Detection efficiencies $\epsilon$ and detection rates $D_{R}$ of the six studied A+, Voyager, and NG networks are plotted as functions of redshift $z$. The circles (squares) denote the values for events with SNR $\rho\geq 10$ ($\rho\geq 100$). The thick, black lines in the rate panels are the cosmic BNS and BBH merger rates, see Sec. III. The fit lines in the efficiency panels are sigmoid fits with $f_{\rm sigmoid}=\Big{(}\frac{1+b}{1+b\,\mathrm{e}^{ax}}\Big{)}^{c}$. The fit parameters for the curves are given in Appendix C.
@@ III.4 Merger and detection rates
[…]
@@
Figure 3: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BNS mergers observed in the six studied A+, Voyager, and NG networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 4: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BBH mergers observed in the six studied A+, Voyager, and NG networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@ IV.2 Signal visibility
Metric   $\Omega_{90}$ $({\rm deg^{2}})$   $\Delta D_{L}/D_{L}$
Quality   $\leq 1$ $\leq 0.1$ $\leq 0.01$  $\leq 0.1$ $\leq 0.01$
@@ V.1 Three-dimensional localization
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors the inclination angle $\Delta\iota$ for BNS mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BBH mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Metric   $\Omega_{90}$ $({\rm deg^{2}})$   $\Delta D_{L}/D_{L}$
Quality   $\leq 1$ $\leq 0.1$ $\leq 0.01$  $\leq 0.1$ $\leq 0.01$
@@ V.2 Early warning alerts
Figure 9: Cumulative histograms for the SNR $\rho$ and 90%-credible sky area $\Omega_{90}$ for BNS mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$ 2, 5, and 10 minutes before merger. The $\Omega_{90}$ panel is obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@ VI Rare and loud events
Figure 10: Histograms of the SNR $\rho_{{\rm post}\textnormal{-}{\rm merger}}$ in the post-merger signal of the BBH population in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$ and a observation time of 10 years.

### v3

@@ Abstract
In this study, we use simple performance metrics to assess the science capabilities of future ground-based gravitational-wave detector networks—composed of A+ or Voyager upgrades to the LIGO, Virgo, and KAGRA observatories and proposed next generation observatories such as Cosmic Explorer and Einstein Telescope. These metrics refer to coalescences of binary neutron stars (BNSs) and binary black holes (BBHs) and include: (i) network detection efficiency and detection rate of cosmological sources as a function of redshift, (ii) signal-to-noise ratios and the accuracy with which intrinsic and extrinsic parameters would be measured, and (iii) enabling multimessenger astronomy with gravitational waves by accurate 3D localization and early warning alerts. We further discuss the science enabled by the small population of rare and extremely loud events. While imminent upgrades will provide impressive advances in all these metrics, next generation observatories will deliver an improvement of an order-of-magnitude or more in most metrics. In fact, a network containing two or three such facilities will detect half of all the BNS and BBH mergers up to a redshift of $z=1$ and $z=20$, respectively, give access to hundreds of BNSs and ten thousand BBHs with signal-to-noise ratios exceeding 100, readily localize hundreds to thousands of mergers to within $1\,{\rm deg^{2}}$ on the sky and better than 10% in luminosity distance, respectively, and consequently, enable mutlimessenger astronomy through follow-up surveys in the electromagnetic spectrum several times a week. Such networks will further shed light on potential cosmological merger populations and detect an abundance of high-fidelity BNS and BBH signals which will allow investigations of the high-density regime of matter at an unprecedented level and enable precision tests of general relativity in the strong-field regime, respectively.
@@ I.3 Future of gravitational-wave astronomy
There have been extensive studies on the capability of next generation detectors and the scientific discoveries they enable Sathyaprakash:2009xt ; Punturo:2010zza ; Zhao:2010sz ; Sathyaprakash:2012jk ; Vitale:2016icu ; Vitale:2018nif ; Adhikari:2019zpy ; Hall:2019xmm ; Sathyaprakash:2019nnu ; Sathyaprakash:2019rom ; Sathyaprakash:2019yqt ; Kalogera:2019sui ; Maggiore:2019uih ; Evans:2021gyd ; Saleem:2021iwi ; Bailes:2021tot ; Kalogera:2021bya ; Yang:2021qge ; Jin:2022qnj ; Jin:2020hmc ; Jin:2021pcv ; Zhang:2019loq . In this study we will focus in detail on the potential of the Voyager, ET, and CE detector generations using the Voy, ET, and CE noise curves shown in Fig. 1. While out of the scope of our investigation, the triangular geometry proposed for ET facilitates a particularly powerful probe of the detector’s sensitivity and data quality. In a triangular configuration, ET consists of three V-shaped detectors, each with an opening angle of 60 degrees, rotated relative to each other by 120 degrees. The sum of the responses of the three detectors, called the null stream Branchesi:2023mws , is, by construction, devoid of any GW signals Sathyaprakash:2009xt . This null stream enables an unbiased estimation of the detector’s sensitivity, free from the confusion noise from the background of unresolved GW sources, and also provides a way of measuring non-stationary, incoherent noise artifacts Goncharov:2022dgl .
@@ I.4 Organization of the paper
In Sec. II we will introduce the detector networks we are considering in this study, the properties of the source populations we are examining, and the analysis methodology we are using to assess the measurement capabilities of detector networks. In the rest of the paper we will introduce and discuss the metrics used to assess the relative performance of different networks. In Sec. III we will discuss the efficiency of detector networks in detecting signals as a function of redshift (or luminosity distance). Detection efficiency will inform us of the completeness of the observed population with respect to the full underlying population of mergers. Sec. III will also discuss the cosmic merger rate of compact binary coalescences inferred using current LIGO-Virgo observations, a model for the redshift evolution of the star formation rate and metallicity, and the resulting detection rates of the networks as functions of the redshift. In Sec. IV we will present the evaluation of the chosen detector networks and a forecast of the science questions that could be addressed by different generations of networks. Enabling multimessenger astronomy will require accurate 3D sky localization of compact binary coalescences, which will be the focus of Sec. V. This Section will also include forecasts for early warning alerts that would be possible with future detectors, with alerts sent to astronomers 10 minutes, 5 minutes, and 2 minutes before the merger occurs and the corresponding 3D sky localization of the event. Sec. VI is devoted to discussing the science enabled by rare events with SNRs larger than 300 or even 1000, e.g. observing higher multipole modes, testing the black hole no-hair theorem to exquisite precession, etc. Sec. VII provides a brief summary, limitations of the study, and how they might be improved and other studies that would be necessary to firm up the science potential of future detectors.
@@ II.1 Detector networks
Although the choice of when a specific A+ configuration is included in a NG network seems arbitrary, this resulted from our judgement that when a NG detector comes online the corresponding region’s 2G+ detector(s) may no longer be in operation. For example, Virgo is not likely to be operational when ET is built, LIGO-Hanford and LIGO-Livingston may not be operating at the same time as CE. Should it prove necessary to assess the performance of a network comprised of a different combination of detectors, e.g. HLI+C, it would be straightforward to do so with gwbench, see Sec. II.3 for details.
@@ II.2 Source populations
Throughout this paper, we examine two source populations comprising either BNSs or BBHs. We assume the local (i.e. $z=0$) merger rate for BNS and BBH to be $R_{\rm BNS}=320^{+490}_{-240}\,{\rm Gpc^{-3}\,yr^{-1}}$ and $R_{\rm BBH}=24^{+14}_{-7}\,{\rm Gpc^{-3}\,yr^{-1}},$ respectively Abbott:2020khf , as determined by the first and second gravitational-wave transient catalogs LIGOScientific:2018mvr ; Abbott:2020niy . Redshift evolution of the rates is discussed in Sec. III.4.
Each population was uniformly distributed in six redshift bins, $z\in[0.02,0.5],\,[0.5,1],\,[1,2],\,[2,4],\,[4,10],\,[10,50]$, with 250,000 injections per bin. The corresponding luminosity distances $D_{L}$ were calculated with the Planck18 cosmology of astropy. Uniform sampling over several redshift bins provides us with good parameter sampling. However, the probability density functions of the populations are scaled appropriately to capture the redshift-dependent merger rates (see Sec. IV.1 for details).
The injections were further uniformly sampled over sky positions ($\alpha$ and $\cos(\delta)$) and binary orientation angles ($\cos(\iota)$ and $\psi$), with right ascension $\alpha$, declination $\delta$, inclination $\iota$, and polarization angle $\psi$. The injected spins were chosen to be aligned with the orbital angular momentum ($\chi_{1x}=\chi_{1y}=\chi_{2x}=\chi_{2y}=0$) while the $z$-components were uniformly sampled, for BNS $\chi_{1z},\,\chi_{2z}\in[-0.05,0,05]$ and for BBH $\chi_{1z},\,\chi_{2z}\in[-0.75,0,75]$. The BNS masses were chosen to be normally distributed in $[1,2]\,M_{\odot}$ with mean $\mu=1.35\,M_{\odot}$ and standard deviation $\sigma=0.15\,M_{\odot}$. The BBH masses were chosen to follow the POWER+PEAK distribution described in the Second Gravitational-Wave Transient Catalog population paper Abbott:2020gyp , but with the secondary mass sampled uniformly in $[5\,M_{\odot},m_{1}]$. These parameters are summarized in Tab. 2. Note that differences in the simulated populations and network configurations require cautious comparison with other literature.
@@ II.3 Analysis methodology
We assess the performance of the chosen GW networks using a fast, Fisher-information Cutler:1994ys ; Poisson:1995ef ; Balasubramanian:1995bm ; Arun:2004hn based Python package, gwbench, that we developed and made publicly available Borhanian:2020ypi ; Appendix A reports the settings used in gwbench and Sec. VII.2 summarizes the limitations of the Fisher-information formalism. The Fisher information formalism approximates the parameter posteriors to be Gaussian (assuming Gaussian noise), provides an analytic recipe to calculate the covariance matrix of the posteriors, and thus allows the computation of measurement error estimates for the gravitational waveform parameters. The implementation in gwbench incorporates the effect of Earth’s rotation, which is important for signals lasting for longer than about an hour, and numerical differentiation schemes, enabling access to a host of gravitational waveform models in LAL, the LSC Algorithm Library lalsuite —ultimately enabling the use of more sophisticated waveform models and, therefore, more accurate estimates. gwbench packs this in an easy-to-use fashion and further gives access to signal-to-noise ratio (SNR) calculations, various detector locations and noise curves, and basic sampling methods. Alternatively, the gravitational-wave community has developed other Fisher-information based Python packages such as gwfast Iacovelli:2022mbg and gwfish Dupletsa:2022scg that can be used to perform similar analyses.
@@ III.2 Network efficiency
Figure 2: Detection efficiencies $\epsilon$ and detection rates $D_{R}$, see Eqs. (3) and (9), respectively, of the six studied A+, Voyager, and NG networks are plotted as functions of redshift $z$. The efficiency specifies the value at the indicated redshift $z$ while the detection rate corresponds to the integrated rate up to that redshift. The circles (squares) denote the values for events with SNR $\rho\geq 10$ ($\rho\geq 100$). The thick, black lines in the rate panels are the cosmic BNS and BBH merger rates, see Sec. III. The fit lines in the efficiency panels are sigmoid fits with $f_{\rm sigmoid}=\Big{(}\frac{1+b}{1+b\,\mathrm{e}^{ax}}\Big{)}^{c}$. The fit parameters for the curves are given in Appendix C.
@@ III.4 Merger and detection rates
where $\Omega_{M}$ and $\Omega_{\Lambda}$ are the dark matter and dark energy densities and we have assumed a flat Universe in which dark energy is interpreted as a cosmological constant Sahni:1999gb . The delay time is not the same for all binaries and depends on the configuration of the compact binary when it forms and is modelled by a probability density $P(t_{d}).$ The merger rate density $\dot{n}(z)$ (normally computed in units of Gpc-3 yr-1) in the source’s frame is obtained by integrating the SFR over all possible time delays,
@@
Figure 3: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BNS mergers observed in the six studied A+, Voyager, and NG networks. The histograms were generated from $\sim 4.7\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 4: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BBH mergers observed in the six studied A+, Voyager, and NG networks. The histograms were generated from $\sim 1.2\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@ IV.2 Signal visibility
Metric AAA     $\Omega_{90}$ $({\rm deg^{2}})$ A     $\Delta D_{L}/D_{L}$  A
Quality     $\leq 1$ $\leq 0.1$ $\leq 0.01$A    $\leq 0.1$ $\leq 0.01$
@@ V.1 Three-dimensional localization
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors the inclination angle $\Delta\iota$ for BNS mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$. The histograms were generated from $\sim 1.2\times 10^{5}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BBH mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$. The histograms were generated from $\sim 1.2\times 10^{4}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Metric     $\Omega_{90}$ $({\rm deg^{2}})$ A     $\Delta D_{L}/D_{L}$  A
Quality     $\leq 1$ $\leq 0.1$ $\leq 0.01$A    $\leq 0.1$ $\leq 0.01$
@@ V.2 Early warning alerts
Figure 9: Cumulative histograms for the SNR $\rho$ and 90%-credible sky area $\Omega_{90}$ for BNS mergers observed in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$ 2, 5, and 10 minutes before merger. The histograms were generated from $\sim 1.2\times 10^{5}$ injections sampled according to Sec. IV.1. The $\Omega_{90}$ panel is obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@ VI Rare and loud events
Figure 10: Histograms of the SNR $\rho_{{\rm post}\textnormal{-}{\rm merger}}$ in the post-merger signal of the BBH population in the six studied A+, Voyager, and NG networks for redshifts $z\leq 0.5$ and a observation time of 10 years. The histograms were generated from $\sim 1.2\times 10^{4}$ injections sampled according to Sec. IV.1.